OpenStudy (anonymous):

If Aragorn can dispatch 5 orcs every 11 seconds and Legolas can dispatch 3 orcs every 9 seconds, then together, how many orcs they can dispatch in how many seconds? A) 2 orcs every 5 seconds B) 8 orcs every 20 seconds C) 26 orcs every 33 seconds D) half an orc every second E) genocide

OpenStudy (xhawaiiangirlx):

To solve this problem, all you need to realize is that the rates at which Aragorn and Legolas slaughter orcs can be represented as fractions and then combined to give their combined rate: 5 orcs every 11 seconds = 5/11 3 orcs every 9 seconds = 3/9 = 1/3 You will notice that we simplified 3/9 to 1/3. It’s always a good idea to simplify fractions whenever possible, as this will save time later. Next, we need to give the fractions like denominators so that we can add them together: Mass slaughter of orcs doesn't count as genocide, does it? Mass slaughter of orcs doesn’t count as genocide, does it? Thus, together they can kill 26 orcs every 33 seconds, and the answer is choice C.

OpenStudy (anonymous):

Thanks!

OpenStudy (decarr432):

answer E is my favorite lol